OpenStudy (nthenic_oftime):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work

OpenStudy (nthenic_oftime):

@OllieT could you help out with this one? its my last one :)

OpenStudy (mazmaz1):

baisically where x equals zero so put in zero for x and solvetill you get your intercepts

OpenStudy (mazmaz1):

this is tricky because you have to factor and sutuff because equation = 0 so you should factor it.

OpenStudy (nthenic_oftime):

x = (12+√24)/12 x = (12-√24)/12 Would these be right ? @mazmaz1

OpenStudy (mazmaz1):

oh God im not sure. could you have somone else look into this

OpenStudy (nthenic_oftime):

yeah thanks any way :)

OpenStudy (anonymous):

yes urs on top were right

OpenStudy (anonymous):

ok, −12±24−−√12=−3±6√3

OpenStudy (anonymous):

so final answer i think would be 0.41